OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places. x^2-5x+5=0 The exact solutions are x= the approximate solutions to 3 decimal places are x=

jhonyy9 (jhonyy9):

delta = 25-20=5 x,y=(5+- 5*(1/2))/2 =(5+-2,41)/2= --- x=2,59/2=1,295 --- y=7,41/2=3,705